# Tokenphormer: Structure-aware Multi-token Graph Transformer for Node Classification

### v1

@@ 2.1 Existing Graph Learning Architectures
Figure 2: Tokenphormer generates diverse tokens with different levels of granularity for the target node (red node), respectively walk-tokens (yellow area), SGPM-token (blue area) and hop-token (green area), comprehensively mining essential information from the whole graph. Then, all these tokens are constructed into a sequence and fed into the Transformer-based backbone to jointly learn the final node representation. Finally, an MLP-based module is employed for node classification tasks.
@@ 2.2 Token-based Representations
NLP has advanced significantly due to deep learning and large language models (LLMs) like GPT (Radford et al. 2018). These improvements hinge on sophisticated token representations that capture semantic and syntactic nuances, enabling applications like sentiment analysis and machine translation. Techniques like Word2Vec (Mikolov et al. 2013) and GloVe (Pennington, Socher, and Manning 2014) use neural networks to place words in continuous vector spaces, enhancing contextual understanding. Models like GPT (Radford et al. 2018) further improve this with contextualized embeddings that consider surrounding words, leading to nuanced semantic comprehension.
In pursuing advanced token representations, there is a shift towards enhanced tokens within graph nodes (Zhang et al. 2020). Traditional graph Transformers (Yun et al. 2019; Ying et al. 2021) use independent nodes as tokens, resulting in high computational complexity and increased noise. Node2Sequence methods like Gophormer (Zhao et al. 2021) sample ego-graphs and add global nodes to form sequences but still overlook edge information, limiting node representation. The Hop2Token method proposed by NAGphormer (Chen et al. 2023) aggregates each hop into a token but remains coarse-grained. Inspired by subword and character-level tokenization (Mikolov et al. 2013), generating multiple tokens for nodes can better capture graph structure details, nuances, and semantic intricacies, aligning with the trend of comprehensive token representations to enhance graph models.
@@ 3.2 Preliminaries
Markov chain and graph random walk. A Markov chain (Chung 1967) is a stochastic process where transitions between states depend only on the current state, reflecting its memoryless property. It has the recurrence property if it can revisit certain states. A random walk (Lovász 1993) on a graph is a type of Markov chain, where each step depends on the current vertex.
Central limit theorem in graphs. The central limit theorem for Markov chains states that if two independent chains with arbitrary initial distributions share the same transition matrix, their limiting distributions will converge to the same value as time approaches infinity (Jones 2004). The limiting distribution can be reached by a Markov chain with recurrence and aperiodic property, satisfying:
@@ 3.3 Graph Serialization
Case 1: Serializing a graph by a long walk. Leveraging the principles of the central limit theorem (Lemma 1) in the context of graphs offers a viable strategy for graph serialization. When the length of a random walk on a graph tends towards infinity, the initial distribution’s impact diminishes, leading to a more comprehensive depiction of the underlying graph structure. We name the document generated through this method the Graph Document, defined as follows:
@@
Case 2: Serializing nodes by diverse walks. Despite serializing the whole graph into graph document, it is also important to serialize the node’s neighborhood for node classification tasks. To ensure that all sentences in the graph are strongly connected with the target node, the graph serialization approach can involve multiple diverse random walks initiated from the target node. Moreover, We also demonstrate that the coverage speed of a node’s neighborhood increases exponentially through this neighborhood serialization method in Appendix A.2111Appendix is available at: https://github.com/Dodo-D-Caster/Tokenphormer..
@@ 3.4 Graph Tokenization
(1) Uniform and non-backtracking random walk. Uniform random walk randomly selects the next node from the current node’s neighbors. Additionally, we explore non-backtracking random walk, which exhibits faster convergence to its limiting distribution compared to uniform random walk in most cases (Alon et al. 2007).
@@
(2) Neighborhood jump walk and non-backtracking version. Conventional random walk approaches are limited by walk length. While the information from nearby neighbors is more significant, disregarding distant neighbors is not advisable. Inspired by dilated convolution (Yu and Koltun 2016), which has been proved to be effective (Contreras, Ceberio, and Kreinovich 2021), we define the neighborhood jump walk as a random walk with dilated transition probability. Unlike uniform random walks, the neighborhood jump walk can jump to any node within its $k$-hop neighborhood and continue recursively. Thus, the neighborhood jump walk’s reachable area expands by the neighborhood’s radius. Below is the definition of the neighborhood jump walk:
@@
Self-supervised Graph Pre-trained Model, SGPM. Self-supervised pre-trained models have emerged as a highly effective approach for learning representations from unlabeled NLP data. Inspired by this, we introduce SGPM to learn from graph document unsupervisedly.
(1) Document generation. As stated in Definition 1, a graph document consists of numerous graph sentences, which are walks generated by individual nodes with varying lengths. In the pre-training phase of SGPM, preserving the raw node connections is crucial for learning the graph’s structure and contextual information. Thus, we use non-backtracking random walks as our final choice, as they extend the walk while maintaining original relationships. The walk length follows a normal distribution, $\mathcal{N}(\mu,\sigma^{2})$, where the mean $\mu$ corresponds to the graph’s radius $R$ and the standard deviation $\sigma$ is set to 1.
(2) Tokenizer and input representation. After generating the graph document, the next step involves tokenizing the sentences. Similar to tokenization in NLP, a vocabulary is created to encompass all tokens. In graph structures, each node maps directly to a corresponding token. Additionally, five special tokens are added: PAD for sequence padding, UNK for unknown nodes, CLS and SEP for sequence boundaries, and MASK for replacing hidden nodes.
@@
(3) Graph pre-training process. Based on experience from NLP, we employ the Masked Language Model (MLM) (Devlin et al. 2019) as SGPM’s pre-train task. Let $\mathbf{H}=[h_{1},h_{2},\ldots,h_{N}]$ represent the input sequence of tokens, and $N$ the total number of tokens. Initially, 15% of the tokens are randomly replaced with the special token MASK. This masked sequence is then input into the bidirectional Transformer model, which generates individual scores $\mathbf{S}=[s_{1},s_{2},\ldots,s_{N}]$ for each token. These scores represent the likelihood of each token given its context and are then compared to the actual ground truth values $\mathbf{Y}=[y_{1},y_{2},\ldots,y_{N}]$, where $y_{i}$ is the true label of the masked token $h_{i}$. The cross-entropy loss is calculated as:
@@ 5 Conclusion
This paper proposes Tokenphormer, a novel graph Transformer architecture that addresses the limitations of MPNNs and graph Transformers by incorporating node contextual information and generating expressive tokens for each node. Specifically, to make the tokens more fine-grained, walk-tokens are generated through a diverse set of walks, incorporating four walk types with varying receptive fields to explore graphs flexibly. Moreover, comprehensiveness is ensured through SGPM-token and hop-token. SGPM-tokens are generated through SGPM to capture global information, extending the length limit of walk-tokens. Hop-tokens are generated through decoupling message passing layers to ensure comprehensive coverage of local information, expanding the density limit of walk-tokens. Tokenphormer has demonstrated state-of-the-art performance in node classification tasks across diverse graph datasets, demonstrating its effectiveness in capturing structural information from raw graphs. We believe that Tokenphormer has the potential to be used in a variety of graph downstream tasks.
@@ A.3 Analysis of tokens
SGPM-token. Appendix A.1 has demonstrated that the graph document can distinguish between different graphs, showcasing its proficiency in capturing the global information of the graph. The the SGPM-token, derived through SGPM, is critical in acquiring contextual information for each node within the graph document. With the generation of numerous walks for each node, the graph document ensures comprehensive coverage for every node. Additionally, the length of the graph sentence follows a normal distribution, with a mean equal to the graph’s radius. This intentional design choice aligns the receptive field of the graph sentence, starting from each node, with the global characteristics of the graph. Furthermore, the graph sentences enable the exploration of far nodes by utilizing the non-backtracking random walk. This approach effectively reduces redundancy, allowing nodes to acquire diverse contexts and enriching the semantic information they gather (Chen et al. 2022).
Hop-token. The hop-token comes from the decoupled message passing layer, and its calculation is:
@@
Walk-token. Instances arise where the hop-token falls short in distinguishing between certain situations. This is exemplified in Fig. 6, where (a) and (b) represent graphs with distinct structures, and (a.1) and (b.2) illustrate the message passing trees for node 1 in their respective graphs. In the depicted figure, the horizontal layers can be interpreted as hop-tokens, while the vertical paths represent walk-tokens. It is evident that when considering a 2-hop message passing scenario, hop-tokens cannot differentiate between graphs (a) and (b). However, walk-tokens, exemplified by (a.2) and (b.2), exhibit the potential to distinguish between the two graphs successfully.
@@ A.4 Complexity Analysis
Time complexity. The time complexity of Tokenphormer primarily relies on the self-attention module of the Transformer. For each node in the graph, we perform computations proportional to the square of the number of tokens $N_{t}$ and the dimension of the feature $d_{F}$, thus resulting in a time complexity of $O(|V|\cdot N_{t}^{2}\cdot d_{F})$, where $|V|$ denotes the number of nodes.
Space complexity. The space complexity is determined by the number of model parameters and the outputs of each layer. It can be broken down into two main parts: 1) The Transformer layer contributes $O(d_{F}^{2}\cdot N_{L})$, where $N_{L}$ is the number of Transformer layers. 2) The attention matrix and hidden node representations add $O(S_{b}\cdot N_{t}^{2}+S_{b}\cdot N_{t}\cdot d_{F})$, where $S_{b}$ denotes the batch size. Thus, the total space complexity is $O(d_{F}^{2}\cdot N_{L}+S_{b}\cdot N_{t}^{2}+S_{b}\cdot N_{t}\cdot d_{F})$.
@@ B.1 Experiment Setup
Datasets. To comprehensively evaluate the effectiveness of Tokenphormer, we conduct experiments on six homogeneous graph datasets with varying node numbers, ranging from 2,708 to 19,717. The datasets include small-sized datasets like Cora and Citeseer (around 3,000 nodes), medium-sized datasets like Flickr and Amazon Photo (around 7,500 nodes), and large-sized datasets like DBLP and Pubmed (around 18,000 nodes). We apply a 60%/20%/20% train/val/test split (consistent with Gophormer (Zhao et al. 2021) and NAGphormer (Chen et al. 2023)) for these benchmark datasets. The detailed dataset information is presented in Table 3.
@@
Baselines. To comprehensively assess the effectiveness of Tokenphormer in the context of node classification tasks, a comparative analysis is conducted against a diverse set of 16 advanced baseline models. The evaluated models encompass eight prominent MPNN methods, including GCN (Kipf and Welling 2017), GAT (Veličković et al. 2018), GraphSage (Hamilton, Ying, and Leskovec 2017), APPNP (Klicpera, Bojchevski, and Günnemann 2018), JKNet (Xu et al. 2018), GPR-GNN (Chien et al. 2021), GRAND+ (Feng et al. 2022), GATv2 (Brody, Alon, and Yahav 2021) and eight innovative graph Transformer models, including GT (Dwivedi and Bresson 2021), Graphormer (Ying et al. 2021), SAN (Kreuzer et al. 2021), Gophormer (Zhao et al. 2021), ANS-GT (Zhang et al. 2022), GraphGPS (Rampášek et al. 2022), Exphormer (Shirzad et al. 2023), Gapformer (Liu et al. 2023), and NAGphormer  (Chen et al. 2023).
@@ B.3 Ablation Study
SGPM-token. SGPM-token captures contextual and global information of the selected node. Without SGPM-token, the accuracy of Tokenphormer drops in all six datasets.
Walk-token. Tokenphormer’s accuracy decreases the most when walk-token are omitted. This is because walk-tokens are able to capture fine-grained information from a large neighborhood.
Hop-token. Hop-token ensures coverage in $k$-hop neighborhoods. However, for most datasets, large number of walk-tokens already cover all nodes in $k$-hop neighborhoods, resulting in a smaller accuracy decrease when hop-token is withdrawn.

### v2

@@ 2.1 Existing Graph Learning Architectures
Figure 2: Framework. RW refers to random walk while NJW stands for neighborhood jump walk. Tokenphormer generates diverse tokens with different levels of granularity for the target node (red node), respectively walk-tokens (yellow area), SGPM-token (blue area) and hop-token (green area), comprehensively mining essential information from the whole graph. Then, all these tokens are constructed into a sequence and fed into the Transformer-based backbone to jointly learn the final node representation. Finally, an MLP-based module is employed for node classification tasks.
@@ 2.2 Token-based Representations
NLP has advanced significantly due to deep learning and large language models (LLMs) like GPT (Radford et al. 2018). These improvements hinge on sophisticated token representations that capture semantic and syntactic nuances, enabling applications like sentiment analysis and machine translation. Techniques like Word2Vec (Mikolov et al. 2013) and GloVe (Pennington, Socher, and Manning 2014) use neural networks to place words in continuous vector spaces, enhancing contextual understanding. Models like BERT (Devlin et al. 2019) and GPT (Radford et al. 2018) further improve this with contextualized embeddings that consider surrounding words, leading to nuanced semantic comprehension.
In pursuing advanced token representations, there is a shift towards enhanced tokens within graph nodes (Zhang et al. 2020). Traditional graph Transformers (Yun et al. 2019; Ying et al. 2021) use independent nodes as tokens, resulting in high computational complexity and increased noise. Node2Sequence methods like Gophormer (Zhao et al. 2021) sample ego-graphs and add global nodes to form sequences but still overlook edge information, limiting node representation. The Hop2Token method proposed by NAGphormer (Chen et al. 2023) aggregates each hop into a token but remains coarse-grained. To make the input tokens of the graph Transformers more fine-grained, inspired by subword and character-level tokenization (Mikolov et al. 2013), generating multiple tokens for nodes can better capture graph structure details, nuances, and semantic intricacies, aligning with the trend of comprehensive token representations to enhance graph models.
@@ 3.2 Preliminaries
Markov chain and graph random walk.
A Markov chain (Chung 1967) is a stochastic process where transitions between states depend only on the current state, reflecting its memoryless property. It has the recurrence property if it can revisit certain states. A random walk (Lovász 1993) on a graph is a type of Markov chain, where each step depends on the current vertex.
Central limit theorem in graphs.
The central limit theorem for Markov chains states that if two independent chains with arbitrary initial distributions share the same transition matrix, their limiting distributions will converge to the same value as time approaches infinity (Jones 2004).
The limiting distribution can be reached by a Markov chain with recurrence and aperiodic property, satisfying:
@@ 3.3 Graph Serialization
Case 1: Serializing a graph by a long walk.
Leveraging the principles of the central limit theorem (Lemma 1) in the context of graphs offers a viable strategy for graph serialization. When the length of a random walk on a graph tends towards infinity, the initial distribution’s impact diminishes, leading to a more comprehensive depiction of the underlying graph structure. We name the document generated through this method the Graph Document, defined as follows:
@@
Case 2: Serializing nodes by diverse walks.
Despite serializing the whole graph into graph document, it is also important to serialize the node’s neighborhood for node classification tasks. To ensure that all sentences in the graph are strongly connected with the target node, the graph serialization approach can involve multiple diverse random walks initiated from the target node. Moreover, We also demonstrate that the coverage speed of a node’s neighborhood increases exponentially through this neighborhood serialization method in Appendix A.2.
@@ 3.4 Graph Tokenization
(1) Uniform and non-backtracking random walk.
Uniform random walk randomly selects the next node from the current node’s neighbors. Additionally, we explore non-backtracking random walk, which exhibits faster convergence to its limiting distribution compared to uniform random walk in most cases (Alon et al. 2007).
@@
(2) Neighborhood jump walk and non-backtracking version.
Conventional random walk approaches are limited by walk length. While the information from nearby neighbors is more significant, disregarding distant neighbors is not advisable. Inspired by dilated convolution (Yu and Koltun 2016), which has been proved to be effective (Contreras, Ceberio, and Kreinovich 2021), we define the neighborhood jump walk as a random walk with dilated transition probability. Unlike uniform random walks, the neighborhood jump walk can jump to any node within its $k$-hop neighborhood and continue recursively. Thus, the neighborhood jump walk’s reachable area expands by the neighborhood’s radius. Below is the definition of the neighborhood jump walk:
@@
Self-supervised Graph Pre-trained Model, SGPM.
Self-supervised pre-trained models have emerged as a highly effective approach for learning representations from unlabeled NLP data. Inspired by this, we introduce SGPM to learn from graph document unsupervisedly.
(1) Document generation.
As stated in Definition 1, a graph document consists of numerous graph sentences, which are walks generated by individual nodes with varying lengths. In the pre-training phase of SGPM, preserving the raw node connections is crucial for learning the graph’s structure and contextual information. Thus, we use non-backtracking random walks as our final choice, as they extend the walk while maintaining original relationships. The walk length follows a normal distribution, $\mathcal{N}(\mu,\sigma^{2})$, where the mean $\mu$ corresponds to the graph’s radius $R$ and the standard deviation $\sigma$ is set to 1.
(2) Tokenizer and input representation.
After generating the graph document, the next step involves tokenizing the sentences. Similar to tokenization in NLP, a vocabulary is created to encompass all tokens. In graph structures, each node maps directly to a corresponding token. Additionally, five special tokens are added: PAD for sequence padding, UNK for unknown nodes, CLS and SEP for sequence boundaries, and MASK for replacing hidden nodes.
@@
(3) Graph pre-training process.
Based on experience from NLP, we employ the Masked Language Model (MLM) (Devlin et al. 2019) as SGPM’s pre-train task. Let $\mathbf{H}=[h_{1},h_{2},\ldots,h_{N}]$ represent the input sequence of tokens, and $N$ the total number of tokens. Initially, 15% of the tokens are randomly replaced with the special token MASK. This masked sequence is then input into the bidirectional Transformer model, which generates individual scores $\mathbf{S}=[s_{1},s_{2},\ldots,s_{N}]$ for each token. These scores represent the likelihood of each token given its context and are then compared to the actual ground truth values $\mathbf{Y}=[y_{1},y_{2},\ldots,y_{N}]$, where $y_{i}$ is the true label of the masked token $h_{i}$. The cross-entropy loss is calculated as:
@@ 5 Conclusion
This paper proposes Tokenphormer, a novel graph Transformer that overcomes the limitations of MPNNs and graph Transformers by generating fine-grained and comprehensive node tokens. Walk-tokens, created through diverse walks with varying walk tendencies and receptive fields, ensure fine-grained exploration of the graph. SGPM-tokens capture global information, extending the length limit of walk-tokens for greater comprehensiveness, while hop-tokens, generated by decoupling message-passing layers, enhance local coverage. Tokenphormer achieves state-of-the-art performance on node classification on diverse datasets, demonstrating its effectiveness for various graph tasks.
@@ A.3 Analysis of tokens
SGPM-token.
Appendix A.1 has demonstrated that the graph document can distinguish between different graphs, showcasing its proficiency in capturing the global information of the graph. The the SGPM-token, derived through SGPM, is critical in acquiring contextual information for each node within the graph document. With the generation of numerous walks for each node, the graph document ensures comprehensive coverage for every node. Additionally, the length of the graph sentence follows a normal distribution, with a mean equal to the graph’s radius. This intentional design choice aligns the receptive field of the graph sentence, starting from each node, with the global characteristics of the graph. Furthermore, the graph sentences enable the exploration of far nodes by utilizing the non-backtracking random walk. This approach effectively reduces redundancy, allowing nodes to acquire diverse contexts and enriching the semantic information they gather (Chen et al. 2022).
Hop-token.
The hop-token comes from the decoupled message passing layer, and its calculation is:
@@
Walk-token.
Instances arise where the hop-token falls short in distinguishing between certain situations. This is exemplified in Fig. 6, where (a) and (b) represent graphs with distinct structures, and (a.1) and (b.2) illustrate the message passing trees for node 1 in their respective graphs. In the depicted figure, the horizontal layers can be interpreted as hop-tokens, while the vertical paths represent walk-tokens. It is evident that when considering a 2-hop message passing scenario, hop-tokens cannot differentiate between graphs (a) and (b). However, walk-tokens, exemplified by (a.2) and (b.2), exhibit the potential to distinguish between the two graphs successfully.
@@ A.4 Complexity Analysis
Time complexity.
The time complexity of Tokenphormer primarily relies on the self-attention module of the Transformer. For each node in the graph, we perform computations proportional to the square of the number of tokens $N_{t}$ and the dimension of the feature $d_{F}$, thus resulting in a time complexity of $O(|V|\cdot N_{t}^{2}\cdot d_{F})$, where $|V|$ denotes the number of nodes.
Space complexity.
The space complexity is determined by the number of model parameters and the outputs of each layer. It can be broken down into two main parts: 1) The Transformer layer contributes $O(d_{F}^{2}\cdot N_{L})$, where $N_{L}$ is the number of Transformer layers. 2) The attention matrix and hidden node representations add $O(S_{b}\cdot N_{t}^{2}+S_{b}\cdot N_{t}\cdot d_{F})$, where $S_{b}$ denotes the batch size. Thus, the total space complexity is $O(d_{F}^{2}\cdot N_{L}+S_{b}\cdot N_{t}^{2}+S_{b}\cdot N_{t}\cdot d_{F})$.
@@ B.1 Experiment Setup
Datasets.
To comprehensively evaluate the effectiveness of Tokenphormer, we conduct experiments on six homogeneous graph datasets with varying node numbers, ranging from 2,708 to 19,717. The datasets include small-sized datasets like Cora and Citeseer (around 3,000 nodes), medium-sized datasets like Flickr and Amazon Photo (around 7,500 nodes), and large-sized datasets like DBLP and Pubmed (around 18,000 nodes). We apply a 60%/20%/20% train/val/test split (consistent with Gophormer (Zhao et al. 2021) and NAGphormer (Chen et al. 2023)) for these benchmark datasets. The detailed dataset information is presented in Table 3.
@@
Baselines.
To comprehensively assess the effectiveness of Tokenphormer in the context of node classification tasks, a comparative analysis is conducted against a diverse set of 16 advanced baseline models. The evaluated models encompass eight prominent MPNN methods, including GCN (Kipf and Welling 2017), GAT (Veličković et al. 2018), GraphSage (Hamilton, Ying, and Leskovec 2017), APPNP (Klicpera, Bojchevski, and Günnemann 2018), JKNet (Xu et al. 2018), GPR-GNN (Chien et al. 2021), GRAND+ (Feng et al. 2022), GATv2 (Brody, Alon, and Yahav 2021) and eight innovative graph Transformer models, including GT (Dwivedi and Bresson 2021), Graphormer (Ying et al. 2021), SAN (Kreuzer et al. 2021), Gophormer (Zhao et al. 2021), ANS-GT (Zhang et al. 2022), GraphGPS (Rampášek et al. 2022), Exphormer (Shirzad et al. 2023), Gapformer (Liu et al. 2023), and NAGphormer  (Chen et al. 2023).
@@ B.3 Ablation Study
SGPM-token.
SGPM-token captures contextual and global information of the selected node. Without SGPM-token, the accuracy of Tokenphormer drops in all six datasets.
Walk-token.
Tokenphormer’s accuracy decreases the most when walk-token are omitted. This is because walk-tokens are able to capture fine-grained information from a large neighborhood.
Hop-token.
Hop-token ensures coverage in $k$-hop neighborhoods. However, for most datasets, large number of walk-tokens already cover all nodes in $k$-hop neighborhoods, resulting in a smaller accuracy decrease when hop-token is withdrawn.